OpenStudy (jkbo):

Please help me Donna secured a $5000 loan from the bank to be paid in 3 years at a 5% rate. How much interest did Donna pay? A.1500 B.7,500 C.750 D.15,000

OpenStudy (anonymous):

\[5000*\frac{ 5 }{ 100 } = \frac{ 25000 }{ 100 } = 250\] This is the interest in one year To spread it over 3 years you multiply by 3 \[250*3=750\]

OpenStudy (jdoe0001):

\(\bf interest = principal\times rate(\textit{in decimals})\times time(\textit{years})\\\quad \\ balance = balance + interest \)

OpenStudy (jkbo):

Thank you both